MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2007/06/26
OBSERVER: Fukui Maiko Paul
CONDITION: clear
COMMENT: We observed intensively gb18 for MOA249, gb10 for MOA209 and gb14 for MOA 217. Good seeing !! Dark D2399(200sec), D2400(60sec) are OK.
Compressor gas pressure: 1.8 MPa
Expected Observation Time: START 18:22 END 07:19


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB7946-80056040324180
LMC1575-15773132600
SMC-000
ALERT-000
GRB-000
TOTAL---6341626780 (07h26m20s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
121212125912121259201212
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
124611121219121111110


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
111101111111
Field13Field14Field15Field16
1100


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
18:00observingclear
19:00observingclear, gb18 for MOA249 and gb10 for MOA209
20:00observingclear, gb18 for MOA249 and gb10 for MOA209
21:00observingclear, gb18 for MOA249 and gb10 for MOA209
22:00observingclear, gb18 for MOA249 and gb10 for MOA209
23:00observingclear, gb14 for MOA 217
00:00observingclear, gb14 for MOA 217
01:00observingclear, gb14 for MOA 217
02:00observingclear, gb14 for MOA 217
03:00observingclear, gb14 for MOA 217
04:00observingclear, gb14 for MOA 217
05:00observingclear, gb14 for MOA 217
06:00observingclear, gb14 for MOA 217
07:00observingclear


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day